Helium boils at –178 o C and a
pressure of 1900 mmHg. What will the
pressure be if the temperature is raised
to 0 o C and the volume is kept constant?
1
Expert's answer
2020-04-03T11:51:38-0400
"\frac{P_1}{T_1}=\frac{P_2}{T_2}\\now, P_1= 1900\ mm Hg\\T_1=273-178\ K=95\ K\\"

"T_2= 273\ K\\"

Putting all values,

We get ,


"\frac{1900}{95}=\frac{P_2}{273}\\or\\P_2=5460mm\ of Hg"